OpenStudy (anonymous):

Solve log3 27 – x = 5

OpenStudy (anonymous):

Change it to exponential: \[\Large 3^5=27-x\] Solve \[\Large x=27-3^5=-216\]

OpenStudy (anonymous):

? i dont get it

OpenStudy (anonymous):

ooh nevermind